OpenStudy (anonymous):

A kilobyte is 2^10. A megabyte is 2^20 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.

OpenStudy (anonymous):

1mb=2^20b= \[\frac{ 2^{20} }{ 2^{10} }=2^{20}*2^{-10}=2^{20-10}=2^{10}kilobytes\]

OpenStudy (anonymous):

so how many kilobytes are in a megabyte and how would i do this?

OpenStudy (phi):

surj gave you the answer in exponential form (just as the question asks for) as an aside, 2^10 is 1024, which is very close to 10^3 (1000)

OpenStudy (anonymous):

So what should i put as the answer its an essay question. could you walk me through each step

OpenStudy (phi):

give your answer in exponential form.

OpenStudy (anonymous):

What surj said, could i use that as the answer

OpenStudy (anonymous):

how would i do that?

OpenStudy (phi):

yes, but you should understand what he did. the question says "Use the exponent rule for division " that rule is "top exponent minus bottom exponent" so 2^20 divided by 2^10 means 2^(20-10) (subtract exponents) to get 2^10

OpenStudy (anonymous):

so if its multiplication you add the exponents and if its division you would subtract?

OpenStudy (phi):

yes, that is exactly the rule.

OpenStudy (phi):

notice what a great rule it is: start with \[ \frac{ 1}{x} \] now use the idea that x to the 0 power is 1, so this is the same as \[ \frac{x^0}{x^1} \] use the subtract rule to get \[ \frac{x^0}{x^1} = x^{0-1} = x^{-1}\] that means \[ \frac{ 1}{x} = x^{-1}\]